### v1

# Prompt Mixing in Diffusion Models using the Black Scholes Algorithm

Abstract
We introduce a novel approach for prompt mixing, aiming to generate images at the intersection of multiple text prompts using pre-trained text-to-image diffusion models. At each time step during diffusion denoising, our algorithm forecasts predictions w.r.t. the generated image and makes informed text conditioning decisions. To do so, we leverage the connection between diffusion models (rooted in non-equilibrium thermodynamics) and the Black-Scholes model for pricing options in Finance, and draw analogies between the variables in both contexts to derive an appropriate algorithm for prompt mixing using the Black Scholes model. Specifically, the parallels between diffusion models and the Black-Scholes model enable us to leverage properties related to the dynamics of the Markovian model derived in the Black-Scholes algorithm. Our prompt-mixing algorithm is data-efficient, meaning it does not need additional training. Furthermore, it operates without human intervention or hyperparameter tuning. We highlight the benefits of our approach by comparing it qualitatively and quantitatively to other prompt mixing techniques, including linear interpolation, alternating prompts, step-wise prompt switching, and CLIP-guided prompt selection across various scenarios such as single object per text prompt, multiple objects per text prompt and objects against backgrounds. Code is available at https://github.com/divyakraman/BlackScholesDiffusion2024.
Figure 1: Our method’s results (Black Scholes, last column) are presented alongside comparisons to prior work. Vanilla stable diffusion (SD) struggles to capture clear characteristics of individual text prompts (notably missing distinct features such as those of the parrot, cat, dog/penguin, and sunset/penguin). Linear interpolation performs poorly due to non-linear manifolds. Alternating sampling and step-wise switching yield low-quality results with artifacts, primarily because they lack intelligent prompt selection during denoising steps (missing characteristics of pizza, artifacts in cat/muffin and dog/penguin mixing, sunset/northern lights not well captured). CLIP-min exhibits bias issues by not modeling diffusion denoising dynamics and prompt selection effectively, which hinders fore-sighted decision making, the generated images are biased towards one of the text prompts. In contrast, our Black-Scholes model generates realistic images that meticulously balance and preserve the characteristics of each individual text prompt. The images are from set 1, set 2, set 3 and set 4 (Refer Section 5.1) respectively.
1 Introduction
Text-to-image diffusion models [13, 14] such as Imagen [33] by Google, Firefly [2] by Adobe, DALLE [7] by Microsoft, and Stable Diffusion [31, 11] by Stability AI can produce photorealistic images from textual descriptions and are typically trained on vast amounts of data using substantial computational resources. The latest versions of these models (as of May 2024) have seen significant improvements in comparison to their initial versions, attributed to various factors, such as increased training data, enhanced text understanding networks, improved image generation networks, and optimized strategies. Data efficient solutions [36], leveraging pre-trained diffusion models, can enhance results without requiring additional data collection. Prompt engineering [37] is one effective approach, involving techniques such as rephrasing prompts to improve model generalization. Another method is using large language models (LLMs) [23, 38] to parse complex prompts and determine useful priors for image generation. Attention maps [9] and prompt mixing techniques [27, 5] also contribute to better results.
Prompt mixing becomes particularly valuable when aiming to generate images that blend different text-image concepts i.e. when aiming to generate images at the intersection [27, 30, 28] of different text-image concepts. For instance, consider generating an image resembling a hybrid of a pink cat and a dog. Current diffusion models [2, 33, 7, 11] excel at straightforwardly combining individual text prompts to create visually appealing combinations. However, these capabilities are often restricted to advanced model versions (version 3) and rely on large-data techniques.
Data efficient prompt mixing techniques can enhance image generation without additional data. While linear interpolation [17] of text embeddings from individual prompts is a simple approach, it may not be optimal due to the highly non-linear nature of the text-image manifold and potential bias issues. Alternatives include switching between prompts during diffusion denoising, either by alternating [18] or using a step-wise technique [27]. However, these methods often require human involvement and careful prompt engineering for optimal results.
The question arises: what’s the best way to switch between text prompts for prompt-mixing? An automated approach considers the model’s varying capabilities with different prompts. While attention maps [15, 9] and layout guidance![42, 10] methods excel at guiding the model toward distinct scene entities, they may not be as effective for blending concepts within the same entity. To address this, during each step of diffusion denoising, the network should automatically focus on the text prompt that corresponds to the image’s deficiencies.
Main contributions.
In this paper, we introduce a novel approach to prompt mixing by integrating concepts from economics and finance. Our objective is to leverage pretrained diffusion models for prompt-mixing to generate images at the intersection of various text-image manifolds. During each step of the diffusion denoising process, our method dynamically conditions on the text prompt that requires the highest level of attention. To achieve this, our algorithm assesses the ‘cost’ associated with each relevant text prompt and selects the optimal conditions for the subsequent step of image generation based on the ‘cost’ that requires maximum optimization.
Key Insight: The inspiration for our approach lies in non-equilibrium thermodynamics [35], which is at the core of the development of diffusion models. These models share a conceptual foundation with the Black Scholes model[26], a Nobel Prize-winning mathematical framework extensively employed in financial markets for pricing European call options[8]. The denoising steps within a diffusion model, aimed at image generation, constitute a Markovian time-series. In our analogy, the image itself represents a valuable ‘stock’ or ‘asset’ that we aim to ‘purchase’ (or generate) at the most favorable ‘cost’ (or alignment with the text prompts). During each step of the diffusion denoising process, we extrapolate the generate latents to compute the ‘stock’ prices at the corresponding time-step, and use the properties of diffusion models to compute the various variables involved in the Black Scholes algorithm. Consequently, we leverage the Black Scholes model to predict a score for each text prompt. This score serves as an indicator of how the image should be conditioned in the subsequent timestep. By adopting this approach, our model dynamically focuses on aspects that require attention, ultimately generating an image that optimally satisfies all relevant text prompts.
We perform experiments on prompts with varying complexities to assess our method’s ability to seamlessly blend different objects and backgrounds across scenarios. Through qualitative and quantitative comparisons against several baselines-such as linear interpolation, prompt switching, and CLIP score-we highlight the superiority of our approach.
2 Related work
Prompt mixing.
Prompt mixing is a technique where different text prompts are used at various steps during the diffusion denoising process. Patashnik et al. [27] follow a step-wise approach, while Aerial Diffusion [18] alternates between prompts to create semantically consistent aerial-view images. Tools such as Image Mixer Diffusion [28] and CLIP Guided Image Mixing [30] follow similar techniques of prompt switching to blend multiple text prompts. However, these methods often require complex hyperparameter tuning.
Mixing step.
The mixing time or mixing step [22] of a Markov chain refers to the duration it takes for the chain to reach its steady-state distribution. Mixing time has found applications in image editing [46] and synthesizing out-of-distribution (OOD) images [45]. When switching between prompts using mixing time, a step-wise approach is followed, akin to the work by Patashnik et al. [27]. The advantage of mixing-time approaches over the method proposed by Patashnik et al. [27] is that it mathematically determines the optimal switching time without requiring complex hyperparameter tuning. To approximate the mixing time, it estimates the radius of the latent space. However, there are a few limitations to these methods. Firstly, most of these approaches are applied to diffusion models trained on relatively small, problem-specific datasets. Secondly, identifying clear boundaries on large-scale foundation models remains an unsolved challenge. Additionally, these methods lack the flexibility to choose the most optimal prompt at each timestep.
Understanding the latent space.
Tangential to our work, there has been substantial work [16, 1, 12] on exploring the latent space. These investigations have yielded valuable insights and practical solutions for downstream tasks such as image editing and manipulation applications [44, 34, 21, 43, 29]. Furthermore, a deeper understanding of the latent space within diffusion models has led to advancements in various methods [31, 21, 40].
3 Background
3.1 The Black Scholes Model
In this section, we provide a brief overview of the Black Scholes pricing model [26] used to determine the price of European call options of assets. which is used to determine the price of European call options on assets. In simple terms, a European call option allows an investor to lock in the price of an asset at any time but permits stock purchase (if desired) only upon expiration. Regardless of whether the stock price moves favorably or unfavorably over time, this option structure remains consistent. Investors rely on the Black-Scholes model to predict stock prices over time and make informed decisions about the optimal timing for stock purchases. The Black Scholes formula involves $5$ key variables:
  1. 1.
    Underlying stock price or spot price $S$: This represents the current price of the asset.
  2. 2.
    Strike price $K$: The strike price is the cost of the asset at the time of expiry.
  3. 3.
    Time to expiration $t$: It measures the time difference between the current moment and the expiry time.
  4. 4.
    Volatility $\sigma$: Volatility reflects the variation in prices of the asset.
  5. 5.
    Risk free rate $r$: The risk-free rate is the minimum return on an investment when the investor faces zero risks.
To obtain the Black Scholes score of purchasing an asset, the spot price $S$ is first multiplied by the standard normal probability distribution function. From this result, to obtain the final cost $C$, the strike price $K$ multiplied by the cumulative standard distribution function is subtracted. Mathematically,
$SN(d_{1})-Ke^{-rt}N(d_{2}),$ (1)
where
$d_{1}=\frac{log\frac{S}{K}+(r+\frac{\sigma^{2}}{2})(t)}{\sigma\sqrt{t}},d_{2}=d_{1}-\sigma\sqrt{t}.$ (2)
Assumptions made by the Black Scholes Model.
The Black Scholes model makes five assumptions: (i) No dividends are paid out during the life of an option, (ii) Market movements are somewhat random, (iii) There are no transaction costs in buying the asset, (iv) The volatility and risk free rate of the underlying asset are known and constant, (v) The returns of the underlying asset are normally distributed.
3.2 Relation to Diffusion Models
The main concept behind diffusion models [13] involves iteratively adding small amounts of random Gaussian noise to transform an initial photorealistic image $x_{0}$ to noise $x_{T}\sim\mathcal{N}(0,I)$ in $T$ steps. This process is known as the forward process. The converse of the theorem states that by starting from random noise $x_{T}\sim\mathcal{N}(0,I)$ and refining it iteratively for $T$ steps, we can generate a photorealistic image $x_{0}$. Since diffusion is a gradual process, the value of $T$ is typically large. At each intermediate timestep, $t\in\{0....T\}$, $x_{t}$ satisfies $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$. $0=\alpha_{T}<\alpha_{T-1}....<\alpha_{1}<\alpha_{0}=1$ are the hyperparameters of the diffusion schedule [13] and $\epsilon_{T}\sim\mathcal{N}(0,I)$. To obtain $x_{t-1}$ at each refinement step, the neural network $f_{\theta}(x_{t},t)$ is applied along with the corresponding random Gaussian noise perturbation. In other words, during each step of diffusion denoising, the known variance in added noise follows a Gaussian distribution [13]. In finance, the volatility refers to the standard deviation in the way stock prices change from one time instant to the next. The mathematical description of diffusion models, derived from non-equilibrium thermodynamics  [35], shares similarities with the derivation of the Black-Scholes model [26] used in pricing European call options within financial markets. Both models emerge from similar assumptions and conditions, underpinned by a shared mathematical structure. Assumptions (i) and (iii) made by the Black Scholes model, which talk about dividends and transaction costs, are irrelevant to diffusion models.
Figure 2: We present more results of our method along with comparisons. Vanilla SD fails to capture clear characteristics of individual text prompts, omitting distinct features such as those related to avocado/raccoon, muffin, parrot, and oil painting style. Linear interpolation generates images not consistent with the prompts, due to issues with non-linear manifolds. CLIP Min. generates images biased towards one of the prompts. Alt. and Step prompt selection methods suffer from artifacts and are not very successful in blending the characteristics of objects corresponding to the individual text prompts - the avocado/raccoon, muffin/dog are not blended well. In the parrot/dog image, the characteristics of the parrot are missing. Alt. generates artifacts in the Times Square/ oil painting image, while Times Square is not characterized well in the image generated by Step. In contrast, the Black-Scholes model adeptly overcomes these limitations, generating realistic images that meticulously balance and preserve the unique characteristics of each individual text prompt. The images are from set 1, set 2, set 3 and set 4 (Refer Section 5.1) respectively.
4 Prompt Mixing using Black Scholes Model
4.1 Problem Formulation
Consider a set of $N$ text prompts denoted as $P_{1},...P_{N}$ and $T$ diffusion denoising steps. Note that text-to-image diffusion architectures often include a text encoder that maps the text prompt to a joint text-image space. Starting from random noise and conditioned on the embedding in the text-image space, the neural network generates the final image.
Our objective is to generate an image that aligns with multiple text prompts simultaneously i.e. at the intersection of the various text-image manifolds of the diffusion models. We pose this task as a prompt mixing [27] problem. During each step of diffusion denoising, our aim is to select the most relevant text prompt (or the text prompt with respect to which the image requires further refining) for conditioning the model, ensuring the generation of an optimal image satisfying all individual text prompts.
4.2 Method
Selecting the most effective text prompt to condition a text-to-image diffusion model can be approached in various ways. Naively switching [18, 27] between prompts is sub-optimal due to the significant human effort involved. Moreover, it fails to address the critical issue of prioritizing the most relevant prompt in an automatic manner. Alternatively, one can compute the CLIP score at each step and move toward the text prompt with the lowest CLIP score. While this approach is reasonable, it overlooks essential factors related to the dynamics [13] of the diffusion denoising process, which significantly impact the final image generation. Leveraging the Black Scholes model allows us to model these dynamics effectively.
We utilize the CLIP score to quantify the “price” of the generated image, treating it as analogous to a stock. The CLIP score has been widely used to measure text-image alignment, making it a suitable metric for evaluating how well our generated image aligns with each text prompt. Consider diffusion denoising at step $i$ out of total steps $T$. We derive the relevant variables for Black Scholes formula as follows:
  1. 1.
    Underlying stock price $S$: The underlying stock price corresponds to the current value of the asset. Let $z_{t}$ represent the predicted latents at timestep $t$. We denote $z_{0,t}$ as the latents of the final predicted image extrapolated from $z_{t}$. In other words, if the denoising process were to proceed in a straightforward manner following the same direction as computed for $z_{t}$, the latents of the final predicted image would be $z_{0,t}$. The underlying stock price, denoted as $S$, is determined by the CLIP score with respect to text prompt $P_{p}$. This CLIP score can be computed using the image decoded from $z_{0,t}$. To maintain values within the range of 00 to $100$, we multiply the CLIP score by $100$.
  2. 2.
    Strike price $K$: This represents the estimated asset price at expiry, following the last step of diffusion denoising, which predicts the final image. Based on our experience with diffusion models, where we found that a CLIP score of approximately $0.25$ indicates reasonable text-image alignment, we opted for a constant value of $0.25$.
  3. 3.
    Time to expiration $t$: $t=T-i$.
  4. 4.
    Volatility $\sigma$: We compute the volatility as the square root of the variance used by the diffusion denoising scheduler at timestep $i$.
  5. 5.
    Risk-free rate $r$: $r=1/T$.
During each iteration $i$ of diffusion denoising, we calculate the Black-Scholes score $b_{i,p}$ for each text prompt $P_{p}$, $p=1...N$ using the given variables and Equation 1 and Equation 2. Subsequently, during the next step of diffusion denoising, we condition on the text prompt with the lowest Black-Scholes score. This approach ensures that the denoising process prioritizes the text prompt associated with the minimum Black-Scholes score for generating an optimal image consistent with all prompts. Please refer to Algorithm 1 for a step-wise description of the method.
Algorithm 1 Black Scholes prompt mixing in backward diffusion enables the diffusion model generate an optimal image at the intersection of multiple text-image manifolds.
1:Initialize latents to random Gaussian noise. $z_{T}\sim\mathcal{N}(0,I)$; T is the total number of diffusion denoising steps.
2:Use the text encoder $\epsilon$ to compute the CLIP embeddings of a (linguistic) combination of the text prompts $\{P\}$. $e\leftarrow\epsilon_{p}(\cup\{P\})$.
3:Initialize Black Scholes variables, strike price $K=0.25\times 100$, $rater=1/T$.
4:// Diffusion denoising - image prediction/generation loop
5:for $t\leftarrow T$ to 00 do
6:     $z_{t-1}=z_{t}-f(z_{t},t,e)$; f is the diffusion UNet.
7:     Use $z_{t-1}$ to compute $z_{0,t-1}$.
8:     Variance $\sigma$ at step $t-1$leftarrow computed using the scheduler of the diffusion model.
9:     for $i\leftarrow 1$ to $N$ do
10:         Spot price $S$leftarrow CLIP score With respect to text prompt $P_{i}$
11:         Time to expiration $\leftarrow t$
12:         Black Scholes score $b_{t,i}$ with respect to prompt $P_{i}$, at timestep $t$, $\leftarrow$ use Equation 1,2.
13:     end for
14:     $P_{min}$leftarrow Text prompt corresponding to min$\{B_{t,i}\}$, $i=1...N$
15:     $e$leftarrow $\epsilon_{p}(P_{min})$
16:end for
5 Experiments and Results
Metrics.
In line with previous research [9], we assess performance using the CLIP Score. We use two variants of the CLIP Score: (i) CLIP-combined, which computes the score by comparing the generated image against a combination of individual text prompts; and (ii) CLIP-add, which averages the CLIP scores for the generated image across each individual text prompt. To compute our metrics, we generate five images for each text prompt and use them all in the computation of our metrics.
Baselines.
Our study examines several baselines: (i) Vanilla SD: Prompt engineering, we use the vanilla stable diffusion model and condition it on a single text prompt effectively describing all individual text prompts, (ii) Lin. Interp.: Direct combination of text embeddings, achieved through linear interpolation between text embeddings [17]. This method equally weights the text embeddings associated with each text prompt. (iii) Switching between text prompts. We consider two variations here: (iii-a) Step: Following Patashnik et. al. [27], we use one text prompt for the initial 7th to 17th denoising steps and then switch to the other text prompt for the remaining steps, (iii-b) Alt.: Following Kothandaraman et. al. [18], we alternate between the two text prompts. (iv) CLIP Min.: Score-based combination, denoted as CLIP-min, where we select the text prompt corresponding to the lowest CLIP score from the previous denoising iteration.
Backbone architecture.
We use the Stable Diffusion 2.1 backbone model in all our experiments. There is no training involved, we use the pretrained model to directly perform inference.
Method CLIP-combined CLIP-add
Vanilla $0.3463$ $0.3012$
Linear Int. $0.2885$ $0.2778$
Alt. Samp. $0.3445$ $0.3098$
CLIP Min. $0.3195$ $0.2955$
Step $0.322$ $0.2997$
Black Scholes $0.3469$ $0.3112$
Table 1: We follow prior work [9, 3] and analyze using CLIP Scores. The Black Scholes algorithm for prompt mixing in diffusion models achieves superior results, compared to other prompt-mixing techniques, as also evidenced by the qualitative results.
5.1 Experimental Settings
We consider the following experimental settings:
Set 1:
Simple text prompts with single objects. In this experiment, we validate our approach using straightforward scenarios, we use two text prompts, each describing one class. Specifically, we consider $17$ classes: [‘a rock’, ‘a coffee mug’, ‘a cute dog’, ‘a pink cat’, ‘a teddy bear’, ‘a robot’, ‘an alien’, ‘an avocado’, ‘a cute raccoon’, ‘a corgi dog’, ‘a parrot’, ‘a car’, ‘a squirrel’, ‘a cute rabbit’, ‘pizza’, ‘muffin’, ‘icecream’], and construct $13617c_{2}=$ 136 text prompts.
Set 2:
Multiple objects per text prompt. To add complexity, blend multiple objects, in the presence of additional objects in the scene. The goal of this experiment is to assess the models’ capabilities in blending the right objects in the scene, and preserving the characteristics of the other objects. We work with three classes of objects: [‘a basket’, ‘a teapot’], [‘apples’, ‘bananas’, ‘a cute cat’, ‘a cute dog’, ‘muffins’], [‘a table’, ‘a carpet’, ‘a bed’] and construct combinations using the first and second set and the second and third set. Additionally, we include the data point “a cat in a bathtub” and “a corgi dog in a bathtub,” resulting in a total of $71$ prompts.
Set 3:
Object actions against backgrounds (prompt mixing w.r.t. object). We investigate the model’s ability to morph objects while considering object-level action information and scene or background context. We consider the following backgrounds, [‘walking in Times Square’, ‘skiing in Times Square’, ‘walking in a beautiful garden’, ‘surfing on the beach’, ‘eating watermelon on the beach’, ‘sitting on a sofa on the beach’, ‘watching northern lights’, ‘watching sunset at a beach’, ‘admiring the opera house in Sydney’, ‘sleeping in a cozy bedroom’, ‘admiring a beautiful waterfall in a forest’, ‘walking in a cherry blossom garden’, ‘walking in a colorful autumn forest’, ‘flying in the sky at sunset’]. The objects are [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. In total, we create $210$ prompts.
Set 4:
Object performing action against a background (prompt mixing w.r.t. background). In this experiment, we explore how the model blends global information and backgrounds while considering objects in various scenarios. We consider the objects [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. The backgrounds and style prompts are [‘walking in Times Square’, ‘sunset’, ‘van gogh style’, ‘oil painting’, ‘a garden with tulips’, ‘northern lights’]. Overall, we have $90$ prompts in this experiment.
5.2 Analysis and Comparisons
We show qualitative results in Figure 1 and Figure 2. More qualitative results can be found in the appendix. The quantitative comparisons are in Figure 1. Our quantitative results are averaged across the metrics for each of the four sets, with $680$, $355$, $1050$, $450$ images respectively. A detailed analysis of our method and benefits over prior art is as follows:
  • Vanilla stable diffusion (SD): Vanilla stable diffusion v2.1 can create plausible combinations of the provided text prompts. However, due to hallucination issues, it is constrained by the distributions it learned during training. As a result, it struggles to capture clear characteristics of individual text prompts. For example, in Figure 1, the distinct features of the parrot, cat, dog/penguin, and sunset/penguin are missing. Similarly, in Figure 2, the avocado/raccoon, muffin, parrot, and oil painting style lack distinct characteristics. Despite this, the generated images align with the broad text prompt descriptions, resulting in a considerable CLIP score.
  • Linear interpolation [17]: Linear interpolation is ill-suited for generating an image that combines two text prompts. This limitation arises from the non-linear nature of the text-image space. While linear interpolation assumes a simple linear relationship, the actual mapping between textual descriptions and image features is intricate. Furthermore, linear interpolation can only traverse a straight path between two points in the latent space, failing to capture nuanced variations or produce novel features beyond the endpoints.
  • Alternating Sampling [18],: Alternating sampling generates features related to both text prompts in the final output image. However, the objects produced suffer from poor definition, unrealistic appearance, and low quality in many instances. Artifacts are prevalent, and several generated images appear implausible. The underlying issue lies in the algorithm’s routine alternation between the two text prompts during diffusion denoising, without adequately emphasizing the most relevant prompt at each step. Despite these limitations, the model achieves a reasonable CLIP Score by broadly aligning with the overall image description, akin to vanilla denoising.
  • Step-wise switching [18]: Step-wise switching is similarly ineffective for the same reasons as Alternating sampling.
  • CLIP-min: The CLIP-min results exhibit bias toward one of the text prompts in most cases, preventing the model from generating an image that aligns with the intersection of both prompts. This bias arises because the model overlooks the dynamics of diffusion denoising and instead selects the text prompt with the lowest CLIP score at each step, resulting in bias-related issues.
  • Black Scholes algorithm: Our approach based on the Black-Scholes algorithm effectively generates realistic images that align with the intersection of the two text prompts. These images exhibit minimal unrealistic artifacts. Additionally, our algorithm successfully preserves individual characteristics corresponding to each text prompt. By modeling the dynamics of diffusion denoising, the algorithm strategically selects the optimal text prompt at each step, achieving a balanced synthesis of both prompts. Notably, our model attains superior quantitative results.
In summary, the Black-Scholes model outperforms all previous methods for prompt mixing by generating realistic images that meticulously preserve and balance the characteristics of each individual text prompt.
Figure 3: Image variations from our method using the Black Scholes model for a single prompt, starting from different random Gaussian noise initializations.
6 Conclusions, Limitations, and Future Work
In this paper, we introduce a novel perspective on prompt mixing, inspired by a probablistic prediction model and its theoretical framework from the domain of Finance. We draw parallels between the Black-Scholes algorithm used in financial markets and diffusion models. From this comparison, we derive an algorithm for prompt mixing, aiming to generate images that align with multiple text prompts. Our extensive analysis showcase both qualitative and quantitative advantages of our approach. Our study has a few limitations, which opens avenues for future work: (i) We follow prior work [9, 3] by using CLIP scores for quantitative evaluation. However, CLIP scores may not be the optimal evaluation strategy, as they can yield similar scores for images with significant quality differences. Due to its learned nature, CLIP is vulnerable to hallucination and bias issues related to its training data. Consequently, it may assign similar scores to images that exhibit general characteristics of the text prompt, without focusing on specific details and features of the scene entities described in the text prompt. To address this, future research could explore more advanced image-language models [25] for evaluation; the development of good metrics for the evaluation of generative models is a widespread issue, and could help better quantify the performance of our method. (ii) Our study focuses on prompt mixing with two prompts. Investigating the impact of using more than two prompts could be a promising direction for future work. (iii) While our method leverages diffusion models, it may not be applicable to non-Gaussian [6] diffusion models or one-step diffusion models [41], as demonstrated in recent research.
Moreover, future work on integrating our approach with recent advancements in attention guidance, layout modeling, and related areas could extend its applicability to critical downstream tasks such as image editing [17, 39, 4], compositionality [24, 3], handling complex prompts [23], text-based view synthesis [18, 19] and personalized image generation [32, 20]. While recognizing the success of recent text-to-image models like Dalle-3, Stable Diffusion 3, Firefly 3, and Imagen 3 in generating images from diverse and complex text prompts, we introduce a data-efficient approach to prompt mixing. Our method offers advantages over previous techniques for prompt-mixing, and future extension could explore integrating our model’s benefits into state-of-the-art text-to-image architectures for data efficient prompt mixing and downstream applications.
Societal impact.
While our method offers valuable tools for generative AI-based content creation, its potential misuse underscores the need for research in watermarking and deepfake detection to mitigate risks.
Acknowledgements.
This work was supported in part by ARO Grants W911NF2110026, W911NF2310046, W911NF2310352 and Army Cooperative Agreement W911NF2120076.